OpenStudy (anonymous):

y'=t-1+y^2 . . . How to solve? Cannot be linear first order ODE due to the y^2, can it?

OpenStudy (anonymous):

t=-y^2+y+1

OpenStudy (anonymous):

Another way of writing it is -t-y^2+y+1 = 0

OpenStudy (anonymous):

Try letting:\[y(t) = t \eta(t) \implies \dot y = \eta(t) + t \dot{\eta}(t) \implies \eta(t)+t \dot \eta (t) = t - 1 + t^2 \eta^2(t)\] From here you could try laplace transforms and get something then convert back?

OpenStudy (anonymous):

I can assume that this is a parabola

OpenStudy (anonymous):

blast234, what are you talking about?

OpenStudy (anonymous):

I think I did it wrong I did not watch out for the ' in the Y

OpenStudy (anonymous):

My bad XD

OpenStudy (anonymous):

In general there is no way to solve non-linear ODEs every time, I just threw out a technique I've seen used.